What is the probability of observing exactly one 3 when rolling a fair die five times?

Question 2 options:

a)

0.0001

b)

0.08

c)

0.40

d)

0.03
Answer #1

This is a binomial distribution question with
n = 5
p = 0.167 (1/6)
q = 1 - p = 0.833
where

Option D is correct
Please hit thumps up if the answer helped you